OpenStudy (anonymous):

Verify identity: sinˆ2(2t)/sinˆ2(t)=4-4sinˆ2(t)

OpenStudy (anonymous):

LHS/RHS proof. \[\huge \frac{\sin^22t}{\sin^2 t}=4-4\sin^2 t\] \[\huge LHS=\frac{\sin^22t}{\sin^2 t}\] \[\huge =\frac{(2\sin t\cos t)^2}{\sin^2 t}\] \[\huge =\frac{4\sin^2 t\cos^2 t}{\sin^2 t}\] \[\huge =4\cos^2 t\] Use your knowledge of trig identities \[\sin^2 t+\cos^2 t=1\] \[\cos^2 t=1-\sin^2 t\] ==================================== \[\huge=4(1-\sin^2 t)\] \[\huge =4-4\sin^2 t\] \[\huge =RHS\]

OpenStudy (anonymous):

Thank you!